Ask
your own question, for FREE!
Statistics
13 Online
In an egg carton there are 12 eggs, of which 9 are hard-boiled and 3 are raw. Six of the eggs are chosen at random to take to a picnic (yes, the draws are made without replacement). Find the chance that at least one of the chosen eggs is raw.
Still Need Help?
Join the QuestionCove community and study together with friends!
find probab. that none of the eggs are raw and subtract it from 1 .
any attempt so far ?
X is a hypergeometric random variable , P ( X > = 1 ) = 1 - P ( X < 1) = 1 - P ( X=0)
|dw:1367062558632:dw|
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
stxandrew1214:
I lwk like a junior need help to talk with her and get to know her but I think sh
thatonenerd:
Use Hvap of water = 40.7 kJ/mol, c of water = 4.184 J/g.u00b0C, and molar mass of HO = 18.
letrasbonitasz:
Typography plays an important role in digital communication. Different font styles can make social media profiles, usernames, messages, and creative designs
Bounty:
tested love: im so over love it acquires patience i don't have trials an tribulations yet my souls the only one taken i give 90 you give 10 you say that you
SaltySuga:
is it possible to tell how deep a cut is by an open wound? like would i be able t
AndreW2002:
I'm trying to change my profile picture but it keeps saying "Either a error occurred, or you tried to upload a gif" but it's not a gif that I'm using, it's
10 hours ago
19 Replies
1 Medal
15 hours ago
5 Replies
0 Medals
2 days ago
3 Replies
1 Medal
14 hours ago
7 Replies
2 Medals
3 days ago
5 Replies
1 Medal
4 days ago
3 Replies
0 Medals